### v1

# A multi-stage Bayesian approach to fit spatial point process models

Abstract
Spatial point process (SPP) models are commonly used to analyze point pattern data, including presence-only data in ecology. Current methods for fitting these models are computationally expensive because they require numerical quadrature and algorithm supervision (i.e., tuning) in the Bayesian setting. We propose a flexible and efficient multi-stage recursive Bayesian approach to fitting SPP models that leverages parallel computing resources to estimate point process model coefficients and derived quantities. We show how this method can be extended to study designs with compact observation windows and allows for posterior prediction of total abundance and points in unobserved areas, which can be used for downstream analyses. We demonstrate this approach using a simulation study and analyze data from aerial imagery surveys to improve our understanding of spatially explicit abundance of harbor seals (Phoca vitulina) in Johns Hopkins Inlet, a protected tidewater glacial fjord in Glacier Bay National Park, Alaska.
1 Introduction
Spatial point process (SPP) models are a class of generative stochastic models that give rise to a random number of irregularly spaced events in a spatial domain. They have been used to analyze point patterns in various fields, such as ecology (e.g., Haase, 1995; Law, 2009; Warton and Shepherd, 2010; Renner et al., 2015), seismology (e.g., Ogata, 1998), and social sciences (e.g., Mohler et al., 2011). One common approach to SPP modeling is to use spatially referenced covariates to fit an inhomogeneous Poisson process (IPP), a point process model in which point intensity varies across space (Gelfand et al., 2010). This approach is especially attractive in applications where point intensity is believed to depend on spatially varying environmental factors.
The stochasticity in an SPP arises from two sources: the random number of point events and their locations, which we refer to as $n$ and $\{\mathbf{s}_{i}\}_{i=1}^{n}$, respectively. A standard IPP likelihood models both data types jointly, allowing spatial heterogeneity and total abundance (i.e., the total number of points in the study domain) to be modeled in a unified framework. This feature is especially useful when the data are partially observed (i.e., total abundance is unknown and may exceed $n$) and inference on total abundance is desired. We extend the IPP likelihood to data observed in compact observation windows, a special case of partially observed data defined by Baddeley et al. (2015). This may arise naturally from certain study designs, such as aerial imagery surveys.
IPP likelihoods are often intensive to compute because they involve an integral over the spatial domain, which is analytically intractable and must be approximated using methods such as numerical quadrature. This is especially inconvenient for Monte Carlo Markov Chain (MCMC; Gelfand and Smith, 1990) methods because the integral is necessarily approximated for each iteration. Furthermore, conventional MCMC algorithms for fitting IPPs often require algorithmic tuning due to a lack of conjugate priors.
One alternative to fitting a Bayesian model using conventional MCMC is to use recursive Bayes approaches, also referred to as Bayesian filtering or sequential inference, to fit the model in stages (Särkkä, 2013). Although these approaches are widely applicable, recursive Bayes approaches for fitting various ecological models have been developed in recent years due to their natural application to hierarchical models and computational efficiency (e.g., Hooten et al., 2023, 2024; Johnson et al., 2022; Leach et al., 2022; McCaslin et al., 2021). These methods often implement a version of the prior-proposal recursive Bayes (PPRB) method described by Hooten et al. (2021). In a two-stage implementation of PPRB, the data are partitioned into two subsets, $\mathbf{y}:=(\mathbf{y}_{1},\mathbf{y}_{2})^{\prime}$. We let $\boldsymbol{\theta}$ represent a vector of model parameters and $\boldsymbol{\theta}]$ represent its prior distribution. We use the bracket notation $\cdot$]’ to represent probability distributions hereafter (Gelfand and Smith, 1990). In the first stage, we use $\mathbf{y}_{1}$ to find a transient posterior distribution $\boldsymbol{\theta}|\mathbf{y}_{1}]\propto[\mathbf{y}_{1}|\boldsymbol{\theta}][\boldsymbol{\theta}]$. Next, in the second stage, we use the transient posterior distribution as a prior and proposal distribution in an MCMC algorithm to obtain a sample from the full posterior distribution
$[\boldsymbol{\theta}|\mathbf{y}]$ $\propto[\mathbf{y}|\boldsymbol{\theta}][\boldsymbol{\theta}]$
$\propto[\mathbf{y}_{2}|\boldsymbol{\theta},\mathbf{y}_{1}][\boldsymbol{\theta}|\mathbf{y}_{1}].$
Critically, using the transient posterior distribution as both the prior and proposal distribution results in cancellations in the Metropolis-Hastings (M-H) ratio for the second stage. Furthermore, the multi-stage implementation allows computationally intensive calculations involved in evaluating the conditional likelihood $\mathbf{y}_{2}|\boldsymbol{\theta},\mathbf{y}_{1}]$ to be computed in parallel in an intermediate stage.
Conventionally, the partitions comprise of data of the same type, but alternative perspectives on partitioning can be helpful in some cases. For instance, PPRB approaches for fitting capture-recapture models partition the data such that the two partitions comprise of the observed binary capture histories and the number of observed individuals, respectively (Hooten et al., 2023, 2024). Similarly, we show that it is natural to partition spatial point pattern data such that the first partition comprises the observed locations $\{\mathbf{s}_{i}\}_{i=1}^{n}$ and the second partition comprises the number of observed point events $n$. We can then fit the data model conditioned on $n$ in the first stage and assimilate $n$ in the second stage. This partitioning scheme allows us to take advantage of well-known strategies that use the conditional data model $\{\mathbf{s}_{i}\}|\boldsymbol{\theta},n]$, such as the logistic and Poisson regression approximations (Aarts et al., 2012; Baddeley et al., 2010; Warton and Shepherd, 2010). We propose a multi-stage method for fitting the IPP using PPRB, along with a variety of strategies for first-stage sampling from the transient posterior distribution, resulting in an efficient and flexible approach to fitting a Bayesian SPP model.
In section 2, we provide an overview for fitting an IPP with spatially explicit covariates. Next, in section 3, we construct a Bayesian IPP model, propose a PPRB procedure for fitting the model, and compare various first-stage sampling strategies. We then demonstrate the multi-stage approach on simulated data in a compact window setting in section 4. Finally, in section 5, we apply our approach to data from aerial imagery surveys to learn spatially explicit abundance of harbor seals in Johns Hopkins Inlet, a protected glacial fjord in Glacier Bay National Park, Alaska.
2 Inhomogeneous Poisson Process
We consider a space of interest, $\mathcal{S}\subset\mathbb{R}^{2}$, and a grid of disjoint subspaces $\{A_{l}\}_{l=1}^{L}$ such that $\mathcal{S}=\cup_{l}A_{l}$. We let $\mathbf{s}_{i}\in\mathcal{S}$ be a two-dimensional vector denoting event location for $i=1,\dots,n$ observed events. We also denote $N(\cdot)$ as a random counting measure and $N$ as the total number of events in $\mathcal{S}$ (i.e., $N=N(\mathcal{S})).$
For an inhomogeneous Poisson process,
  1. i)
    $N(A_{l})$ follows a Poisson distribution with rate parameter $\Lambda(A_{l}):=\int_{A_{l}}\lambda(\mathbf{s})\text{d}\mathbf{s}$.
  2. ii)
    $N(A_{1}),\dots,N(A_{L})$ are conditionally independent random variables for $l=1,\dots,L$.
An IPP can also be thought of as the continuous limit of an independent Poisson count model for an increasingly fine grid of subspaces (i.e., as $L$ approaches $\infty$) over the spatial domain (Fithian and Hastie, 2013). Our method focuses on modeling the continuous intensity function $\lambda$ parametrically using a computationally efficient and flexible Bayesian approach. Additional details and theoretical properties on the IPP are available in Gelfand et al. (2010).
2.1 Complete Likelihood
Assuming we have $p$ covariates associated with each point in $\mathcal{S}$, we can model the IPP intensity as a function of the spatially referenced covariates $\lambda(\mathbf{s})=f(\beta_{0}+\mathbf{x}^{\prime}(\mathbf{s})\boldsymbol{\beta}\bigr{)}$, where $\beta_{0}$ represents the intercept, $\mathbf{x}$ and $\boldsymbol{\beta}$ denote $p\times 1$ vectors of covariates and their corresponding coefficients, respectively, and the function $f$ represents a positive strictly increasing differentiable function. We assume the convention $f(x)=\exp(x)$. If desired, this assumption may be relaxed by using a different link function and/or transforming the covariates using a basis expansion.
The IPP likelihood is then as follows
$\big{[}\{\mathbf{s}_{i}\},n|\beta_{0},\boldsymbol{\beta}\big{]}=\frac{\prod_{i=1}^{n}\lambda(\mathbf{s}_{i})}{n!\exp\bigl{(}\Lambda(\mathbf{\mathcal{S}})\bigr{)}},$ (1)
for $i=1,\dots,n$. We refer to this IPP likelihood that jointly models $\{\mathbf{s}_{i}\}_{i=1}^{n}$ and $n$ as the complete likelihood hereafter.
In practice, the integral $\Lambda(\mathcal{S})$ is not analytically tractable and must be approximated. Using covariate information for grid centers $\mathbf{u}_{l}$ for $l=1,\dots,L$, we can approximate $\Lambda(\mathcal{S})$ with numerical quadrature and use $\{\mathbf{u}_{l}\}_{l=1}^{L}$ as quadrature points, but this can be computationally intensive, especially for large $L$.
2.2 Conditional Likelihood
The IPP likelihood conditional on $n$ is as follows
$\big{[}\{\mathbf{s}_{i}\}|n,\boldsymbol{\beta}\big{]}=\frac{\prod_{i=1}^{n}\lambda(\mathbf{s}_{i})}{\bigl{(}\Lambda(\mathcal{S})\bigr{)}^{n}},$ (2)
for $i=1,\dots,n$, where $\lambda(\mathbf{s})=\exp(\mathbf{x}^{\prime}(\mathbf{s})\boldsymbol{\beta}\bigr{)}$. We refer to this as the conditional likelihood hereafter. It is important to note that unlike in the complete likelihood, $\beta_{0}$ is non-identifiable because it cancels in the numerator and denominator in (2).
The conditional likelihood, sometimes referred to as the conditional IPP, is used in the first stage of our multi-stage algorithm. Additionally, it is used in applications where the number of observed points $n$ is treated as known and fixed a priori. For instance, in telemetric surveys for studying animal movement (Hooten et al., 2017), it is common to fit an SPP model to the observed spatially explicit points for a predetermined number of animal relocations (e.g., Johnson et al., 2013). The conditional likelihood suffers from similar computational challenges in approximating $\Lambda(\mathcal{S})$; however, useful likelihood approximations have been developed to circumvent this issue.
2.2.1 Logistic Regression Approximation
One widely adopted approximation for the conditional likelihood in (2) is the logistic regression approximation, which involves augmenting $m$ background points to the data to represent locations where point events were absent (Baddeley et al., 2010; Warton and Shepherd, 2010; Fithian and Hastie, 2013). Warton and Shepherd (2010) showed that if the number of observed points $n$ is fixed and the number of background points $m$ grows infinitely large, then the logistic regression coefficients (intercept excluded) converge to the maximum likelihood estimates of the IPP coefficients in the conditional likelihood (2).
To implement this approach, we use the Berman-Turner device (Berman and Turner, 1992), where we define an indicator variable, $y(\mathbf{s})=\mathds{1}(\mathbf{s}\in\{\mathbf{s}_{i}\}_{i=1}^{n})$, which indicates whether a location $\mathbf{s}$ is within our set of observed points. By definition, we have $y(\mathbf{s}_{i})=1$ for $i=1,\dots,n$. To obtain points where $y(\mathbf{s})=0$, we generate a uniform random sample of size $m$ within $\mathcal{S}$. For guidance on choosing the number of background points $m$ in practice, see Northrup et al. (2013).
For $m$ background points, we order $y(\mathbf{s}_{i})$ such that $y(\mathbf{s}_{i})=1$ for $i=1,\dots,n$ and $y(\mathbf{s}_{i})=0$ for $i=n+1,\dots,\tilde{n}$ where $\tilde{n}=n+m$. We then fit $y(\mathbf{s}_{i})\sim\text{Bern}(p_{i})$ where $\text{logit}(p_{i})=\beta_{0}+\mathbf{x}^{\prime}(\mathbf{s}_{i})\boldsymbol{\beta}.$ Although the logistic regression intercept is not relevant for the IPP model, note that its inclusion is necessary to adjust for the background sample size (Fieberg et al., 2021).
2.3 Compact Observation Window Extension
[…]
The joint likelihood of $\{\mathbf{s}_{ij}\}$, the set of event locations for each individual $i$ in the $j$th window, and $\{n_{j}\}$, the set of point event counts within each window $j$, becomes
$\big{[}\{\mathbf{s}_{ij}\},\{n_{j}\}|\beta_{0},\boldsymbol{\beta}\big{]}=\frac{\prod_{\{\forall j,n_{j}>0\}}\prod_{i=1}^{n_{j}}\lambda(\mathbf{s}_{ij})}{n!\exp\bigl{(}\Lambda(\cup_{j}\mathcal{S}_{j})\bigr{)}},$ (3)
for $i=1,\dots,n$ and $j=1,\dots,J$ where $n=\sum_{j}n_{j}$. We refer to (3) as the windowed complete likelihood hereafter. The windowed conditional likelihood can be found by conditioning on $\{n_{j}\}$.
3 Bayesian Model
Using the complete likelihood in (1), we can estimate and quantify uncertainty for the parameters of interest, $\beta_{0}$ and $\boldsymbol{\beta}$, using a Bayesian model. The full posterior distribution is as follows:
$\big{[}\beta_{0},\boldsymbol{\beta}|\{\mathbf{s}_{i}\},n\big{]}\propto\big{[}\{\mathbf{s}_{i}\},n|\beta_{0},\boldsymbol{\beta}\big{]}[\beta_{0}][\boldsymbol{\beta}].$ (4)
Recall that, for the complete likelihood in (1), $\Lambda(\mathcal{S})$ is analytically intractable. Using numerical quadrature to approximate this integral often involves computing a large sum that scales with the number of quadrature points. Furthermore, the integral must be computed at each MCMC iteration.
Secondly, we are unable to obtain closed form representations for the full-conditional distributions $\beta_{0}|\cdot]$ and $\boldsymbol{\beta}|\cdot]$. If we use a normal random walk proposal for $\beta_{0}$ and $\boldsymbol{\beta}$, we must perform tuning, which incurs additional computational costs.
Instead of evaluating the full posterior distribution in (4) using a conventional single-stage MCMC algorithm, we propose a multi-stage MCMC algorithm for fitting the complete likelihood that does not require parameter tuning and allows for numerical integration to be computed in parallel.
3.1 Recursive Bayes Algorithm
To construct our multi-stage MCMC algorithm, we apply PPRB following Hooten et al. (2021). Our algorithm follows a similar structure of the two-stage PPRB algorithm summarized in section 1 with the addition of an intermediate stage for parallel computation.
We first show that the complete likelihood in (1) can be decomposed into two terms:
$\big{[}\{\mathbf{s}_{i}\},n|\beta_{0}\ \boldsymbol{\beta}\big{]}$ $=\frac{\prod_{i=1}^{n}\lambda(\mathbf{s}_{i})}{n!\exp\bigl{(}\Lambda(\mathcal{S})\bigr{)}}$ (5)
$=\Biggl{(}\frac{\prod_{i=1}^{n}\lambda(\mathbf{s}_{i})}{\big{(}\Lambda(\mathcal{S})\big{)}^{n}}\Biggr{)}\Biggl{(}\frac{\big{(}\Lambda(\mathcal{S})\big{)}^{n}}{n!\exp\big{(}\Lambda(\mathcal{S})\big{)}}\Biggr{)}$ (6)
$=\big{[}\{\mathbf{s}_{i}\}|\boldsymbol{\beta},n\big{]}[n|\beta_{0},\boldsymbol{\beta}].$ (7)
Moreover, note that the first term in (7) is equivalent to the conditional likelihood in (2). The second term, the likelihood for $n$, follows a Poisson distribution:
$n|\beta_{0},\boldsymbol{\beta}]=\text{Pois}\bigl{(}\Lambda(\mathcal{S})\bigr{)}.$ (8)
The full posterior distribution can then be decomposed as
$\big{[}\beta_{0},\boldsymbol{\beta}|\{\mathbf{s}_{i}\},n\big{]}\propto\big{[}\{\mathbf{s}_{i}\}|\boldsymbol{\beta},n\big{]}[n|\beta_{0},\boldsymbol{\beta}][\beta_{0}][\boldsymbol{\beta}].$ (9)
Evaluating the decomposed posterior distribution in (9) using PPRB results in two stages. In the first stage, we evaluate the transient posterior using the conditional likelihood (2) and the prior for $\boldsymbol{\beta}$
$\big{[}\boldsymbol{\beta}|\{\mathbf{s}_{i}\},n\big{]}\propto\big{[}\{\mathbf{s}_{i}\}|\boldsymbol{\beta},n\big{]}[\boldsymbol{\beta}]$ (10)
to obtain a Bayesian sample, $\big{\{}\boldsymbol{\beta}^{(k)}|\{\mathbf{s}_{i}\},n\big{\}}_{k=1}^{K}$. Recall that in section 2.1.1, we show that fitting the conditional likelihood in (2) can be approximated using logistic regression, which permits a variety of ways to obtain a sample from the transient posterior distribution. We describe and compare first-stage sampling strategies in the context of the multi-stage algorithm in the following section.
In between stages, we use the transient posterior realizations to compute corresponding values of $\Lambda(\mathcal{S})$ in parallel and store them for recall in the following stage. Approximating $\Lambda(\mathcal{S})$ is the most computationally intensive step in evaluating the likelihood for $n$ and parallelization results in significant speedup for the final stage of the algorithm.
In the second stage, we use the transient posterior distribution as a prior and proposal distribution in an MCMC algorithm to account for the stochasticity associated with $n$ in the full posterior distribution. Using the transient posterior distribution (10) as a prior and proposal distribution for $\boldsymbol{\beta}$ results in convenient cancellations involving the data model conditioned on $n$ in the M-H acceptance ratio:
$\alpha$ $=\frac{\big{[}\{\mathbf{s}_{i}\},n|{\beta_{0}}^{(*)},\boldsymbol{\beta}^{(*)}\big{]}\big{[}{\beta_{0}}^{(*)}\big{]}\big{[}{\beta_{0}}^{(k-1)}\big{]}_{*}\big{[}{\boldsymbol{\beta}}^{(*)}\big{]}\big{[}{\boldsymbol{\beta}}^{(k-1)}\big{]}_{*}}{\big{[}\{\mathbf{s}_{i}\},n|{\beta_{0}}^{(k-1)},\boldsymbol{\beta}^{(k-1)}\big{]}\big{[}{\beta_{0}}^{(k-1)}\big{]}\big{[}{\beta_{0}}^{(*)}\big{]}_{*}\big{[}{\boldsymbol{\beta}}^{(k-1)}\big{]}\big{[}{\boldsymbol{\beta}}^{(*)}\big{]}_{*}}$ (11)
$=\frac{\big{[}\{\mathbf{s}_{i}\}|{\beta_{0}}^{(*)},\boldsymbol{\beta}^{(*)}\big{]}\big{[}n|{\beta_{0}}^{(*)},\boldsymbol{\beta}^{(*)}\big{]}\big{[}{\beta_{0}}^{(*)}\big{]}\big{[}{\beta_{0}}^{(k-1)}\big{]}_{*}\big{[}{\boldsymbol{\beta}}^{(*)}\big{]}\big{[}{\boldsymbol{\beta}}^{(k-1)}\big{]}_{*}}{\big{[}\{\mathbf{s}_{i}\}|{\beta_{0}}^{(k-1)},\boldsymbol{\beta}^{(k-1)}\big{]}\big{[}n|{\beta_{0}}^{(k-1)},\boldsymbol{\beta}^{(k-1)}\big{]}\big{[}{\beta_{0}}^{(k-1)}\big{]}\big{[}{\beta_{0}}^{(*)}\big{]}_{*}\big{[}{\boldsymbol{\beta}}^{(k-1)}\big{]}\big{[}{\boldsymbol{\beta}}^{(*)}\big{]}_{*}}$ (12)
$=\frac{\bigl{[}n|{\beta_{0}}^{(*)},\boldsymbol{\beta}^{(*)}\bigr{]}\big{[}{\beta_{0}}^{(*)}\big{]}\big{[}{\beta_{0}}^{(k-1)}\big{]}_{*}}{\bigl{[}n|{\beta_{0}}^{(k-1)},\boldsymbol{\beta}^{(k-1)}\bigr{]}\big{[}{\beta_{0}}^{(k-1)}\big{]}\big{[}{\beta_{0}}^{(*)}\big{]}_{*}}.$ (13)
The ()(*) superscript denotes a proposed value, the (k1)(k-1) superscript denotes the value of a parameter at the (k1)(k-1)th step of the MCMC chain, and $\cdot$$_{*}$ denotes a proposal distribution. Critically, we do not need to perform additional calculations using the point data $\{\mathbf{s}_{i}\}$ at this stage.
Recall that we are unable to learn about $\beta_{0}$ using the first-stage logistic regression approximation. It is, however, crucial to efficiently propose $\beta_{0}$ for the M-H correction in the second stage to obtain reasonable MCMC mixing rates. Conveniently, if we transform the intercept such that $\zeta=e^{\beta_{0}}$, we can obtain a Gibbs sampler using a $\text{Gamma}(a,b)$ prior for $\zeta$. The Gibbs update for $\zeta$ becomes
$\zeta|\cdot]=\text{Gamma}\big{(}a+n,b+\Lambda(\mathcal{S})\big{)}.$ (14)
We then log-transform our posterior sample for $\zeta$ to obtain a sample from the $\beta_{0}$ full-conditional distribution. For a full derivation of the Gibbs update for $\zeta$, see Appendix A.
The Gibbs sampler for $\beta_{0}$ also allows for the M-H ratio in (13) to simplify to
$\alpha=\big{[}n|{\beta_{0}}^{(k)},\boldsymbol{\beta}^{(*)}\big{]}/\big{[}n|{\beta_{0}}^{(k)},\boldsymbol{\beta}^{(k-1)}\big{]}.$ (15)
Thus, each MCMC iteration in the second stage only requires updating $\beta_{0}$ using the Gibbs update in (14) and evaluating $\big{[}n|{\beta_{0}}^{(k)},\boldsymbol{\beta}^{(*)}\big{]}$, both of which are fast using the pre-computed values of $\Lambda(\mathcal{S})$ from the intermediate stage. The full multi-stage algorithm is summarized in Algorithm 1.
First Stage:
 Data: $\{s_{i}\}_{i=1}^{n}$
 Parameters: $\boldsymbol{\beta}$
 for $k=1,\dots,K$ do
      Sample from $\big{[}\boldsymbol{\beta}|\{\mathbf{s}_{i}\},n\big{]}\propto\big{[}\{\mathbf{s}_{i}\}|\boldsymbol{\beta},n\big{]}[\boldsymbol{\beta}].$
   end for
Output: $\big{\{}\boldsymbol{\beta}^{(k)}|\{\mathbf{s}_{i}\},n\big{\}}_{k=1}^{K}$
Intermediate Stage:
 Input: $\big{\{}\boldsymbol{\beta}^{(k)}|\{\mathbf{s}_{i}\},n\big{\}}_{k=1}^{K}$
 do in parallel
    for $k=1,\dots,K$ do
         Approximate $\Lambda^{(k)}(\mathcal{S})$ using first-stage $\boldsymbol{\beta}^{(k)}.$
      end for
   end
Output: $\big{\{}\Lambda^{(k)}(\mathcal{S})\big{\}}_{k=1}^{K}$
Second Stage:
 Data: $n$
 Parameters: $\beta_{0},\,\boldsymbol{\beta}$
   Initialize $\beta_{0}$}, $\boldsymbol{\beta}$} and compute corresponding $\Lambda(\mathcal{S})$.
 for $k=1,\dots,K$ do
      Sample $\zeta^{(k)}\sim\text{Gamma}\big{(}a+n,b+\Lambda^{(k)}(\mathcal{S})\big{)}$.
      Set ${\beta_{0}}^{(k)}=\log\big{(}\zeta^{(k)}\big{)}.$
      Sample uniformly with replacement $\boldsymbol{\beta}^{(*)}\sim\big{\{}\boldsymbol{\beta}^{(k)}|\{\mathbf{s}_{i}\},n\big{\}}_{k=1}^{K}$.
      Fetch corresponding $\Lambda^{(*)}(\mathcal{S})$.
      Set $\boldsymbol{\beta}=\boldsymbol{\beta}^{(*)}$ and $\Lambda(\mathcal{S})=\Lambda^{(*)}(\mathcal{S})$ w.p. $\alpha=\big{[}n|{\beta_{0}}^{(k)},\boldsymbol{\beta}^{(*)}\big{]}/\big{[}n|{\beta_{0}}^{(k)},\boldsymbol{\beta}^{(k-1)}\big{]}$.
      Set $\boldsymbol{\beta}^{(k)}=\boldsymbol{\beta}$ and $\Lambda^{(k)}(\mathcal{S})=\Lambda(\mathcal{S}).$
   end for
Output: $\big{\{}{\beta_{0}}^{(k)},\boldsymbol{\beta}^{(k)}|\{\mathbf{s}_{i}\},n\big{\}}_{k=1}^{K}$
Algorithm 1 Fully Bayesian Multi-stage MCMC
3.2 First-stage Sampling Strategies
One advantage to fitting the Bayesian SPP model in multiple stages is the increased flexibility facilitated by approximating the first stage using a logistic regression model. There are a variety of approaches for fitting a Bayesian logistic regression model such as the Pólya-Gamma data augmentation method described by Polson et al. (2013), which results in a fully Gibbs algorithm and circumvents the need to perform parameter tuning. Other unsupervised Monte Carlo methods for fitting logistic regression can be performed using existing software, such as stan_glm for Hamiltonian Monte Carlo (HMC; Neal, 2011). We refer to fitting the multi-stage procedure with the Pólya-Gamma data augmentation and stan_glm approaches as the Pólya-Gamma (PG) and Bayesian GLM (B-GLM) methods, respectively, in reference to their first-stage sampling strategies.
An alternative to fitting the multi-stage algorithm as fully Bayesian is to fit the first stage using maximum likelihood estimation. This can be attractive because fitting a Bayesian logistic regression model in the first stage can be slow, especially after performing large data augmentations, such as the background points generated from the Berman-Turner device and/or the latent variables required for the Pólya-Gamma Gibbs sampler. This creates a computational bottleneck because the first stage cannot be parallelized and thus does not scale with the number of computational cores. Johnson et al. (2022) proposed a multi-stage algorithm for a different class of models in which the transient posterior distribution is approximated using non-Bayesian likelihood maximization in the first stage. A sample from the approximated transient posterior distribution can then be corrected using an approximate M-H ratio in later stages (Johnson et al., 2022).
Applying a similar approach, we can perform fast maximum likelihood estimation for $\boldsymbol{\beta}$ in the first stage using the glm function in R to fit the logistic regression model described in section 2.2.1. If we assume a flat prior for $\boldsymbol{\beta}$ and standard regularity conditions (i.e., the likelihood is continuously twice-differentiable), we can apply large sample theory to obtain the following approximation:
$\Big{[}\boldsymbol{\beta}|\big{\{}y(\mathbf{s}_{i})\big{\}},n\Big{]}\approx\text{N}\big{(}\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\Sigma}}_{\beta}\big{)},$ (16)
where $\hat{\boldsymbol{\beta}}$ and $\hat{\boldsymbol{\Sigma}}_{\beta}$ are the maximum likelihood estimates and inverse observed Fisher information, respectively. We assume that the sample size is large enough for $\hat{\boldsymbol{\beta}}$ to be identifiable and well-defined. Next, we draw an i.i.d. sample from the approximate Gaussian distribution in (16) as an approximate sample from the transient posterior distribution.
Johnson et al. (2022) showed that the approximate transient posterior distribution in (16) allows for approximate cancellations in the later stage M-H ratio and that parameter estimates obtained using this method are close to those obtained through a conventional single-stage MCMC algorithm. This method can be applied in our multi-stage procedure to obtain an approximate and fast sample from the posterior distribution for $\boldsymbol{\beta}$. Using the Gibbs sampler in (14), the M-H ratio in the second stage again reduces to (15). We refer to this multi-stage method as the GLM approximate (GLM-A) method hereafter.
If more exact estimates are desired, the second-stage M-H correction can be amended to account for the transient posterior approximation in (16). Instead of decomposing the complete likelihood and making approximate cancellations like in (13), we can use the original M-H correction as in (11). Using the Gibbs sampler for $\zeta$ based on the full-conditional distribution in (14) results in the following M-H ratio:
$\alpha=\frac{\big{[}\{\mathbf{s}_{i}\},n|{\beta_{0}}^{(*)},\boldsymbol{\beta}^{(*)}\big{]}\big{[}{\boldsymbol{\beta}}^{(*)}\big{]}\big{[}{\boldsymbol{\beta}}^{(k-1)}\big{]}_{*}}{\big{[}\{\mathbf{s}_{i}\},n|{\beta_{0}}^{(k-1)},\boldsymbol{\beta}^{(k-1)}\big{]}\big{[}{\boldsymbol{\beta}}^{(k-1)}\big{]}\big{[}{\boldsymbol{\beta}}^{(*)}\big{]}_{*}}.$ (17)
To speed up the evaluation of the log-likelihood for $\big{[}\{\mathbf{s}_{i}\},n|{\beta_{0}}^{(*)},\boldsymbol{\beta}^{(*)}\big{]}$, we compute $\sum_{i=1}^{n}\mathbf{x}^{\prime}(\mathbf{s}_{i})\boldsymbol{\beta}$ in parallel in the second stage in addition to $\Lambda(\mathcal{S})$ using the first stage sample. Although the intermediate and second stages involve more calculations compared to the previous multi-stage methods, the overall algorithm results in significant speedup for this model compared to the conventional single-stage method because of the fast first-stage evaluation using maximum likelihood estimation and the ability to parallelize large computations over $k$ in the intermediate stage. We refer to this multi-stage method as the GLM exact (GLM-E) method hereafter.
Both the GLM-A and GLM-E methods result in significant speedup when compared to both the single-stage MCMC algorithm and fully Bayesian multi-stage methods (e.g., PG and B-GLM). Furthermore, the independent samples from the first stage can help prevent sample degeneracy and improve MCMC mixing in the second stage. Table 1 summarizes the strengths of each first-stage sampling strategy.
Property PG B-GLM GLM-E GLM-A
Unsupervised
Fully Bayesian $\times$ $\times$
M-H cancellations $\times$
Exact M-H $\times$
Independent first stage samples $\times$ $\times$
Table 1: Comparison of strengths between various first-stage sampling methods.
3.3 Compact Window Prediction
3.3.1 Posterior Prediction for Total Abundance
In the compact window setting, recall that we do not observe events that occur outside of the compact observation windows. We denote the number of events outside of the observation windows as $n_{0}$. If we assume that the point process across the entire spatial domain is an IPP, then by definition, $n_{0}$ also follows a Poisson distribution:
$\big{[}n_{0}|\beta_{0},\boldsymbol{\beta},\{\mathbf{s}_{ij}\},\{n_{j}\}\big{]}=\text{Pois}\big{(}\Lambda(\mathcal{S}_{0})\big{)}.$ (18)
The posterior predictive distribution for $n_{0}$ is expressed using
$\big{[}n_{0}|\{s_{ij}\},\{n_{j}\}\big{]}=\int\int\big{[}n_{0}|\beta_{0},\boldsymbol{\beta},\{s_{ij}\},\{n_{j}\}\big{]}\big{[}\beta_{0},\boldsymbol{\beta}|\{s_{ij}\},\{n_{j}\}\big{]}\text{d}\beta_{0}\text{d}\boldsymbol{\beta}.$ (19)
After obtaining a posterior sample of $\beta_{0}$ and $\boldsymbol{\beta}$, we obtain a sample from the posterior predictive distribution for $n_{0}$ (19) via composition sampling by drawing ${n_{0}}^{(k)}\sim\text{Pois}\big{(}\Lambda^{(k)}(\mathcal{S}_{0})\big{)}$, where $\Lambda^{(k)}(\mathcal{S}_{0})$ is computed using ${\beta_{0}}^{(k)}$ and $\boldsymbol{\beta}^{(k)}$. Approximating $\Lambda^{(k)}(\mathcal{S}_{0})$ still requires numerical quadrature, but can be computed in parallel for $k=1,\dots,K$. The posterior predictive sample for $n_{0}$ can then be used to perform a finite population correction for the total number of events in the spatial domain, $\mathcal{S}$: $N^{(k)}=n_{0}^{(k)}+n$ for $k=1,\dots,K$, which provides both a point estimate and uncertainty quantification for total abundance.
3.3.2 Posterior Point Simulation
We can simulate posterior predictive event locations in the unobserved region $\mathcal{S}_{0}$ after obtaining estimates for $\lambda(\mathbf{u})$ $\forall\mathbf{u}:\mathbf{u}\in\mathcal{S}_{0}$ using the Lewis-Schedler method for simulating from an IPP (Lewis and Schedler 1979). Alternatively, a new point realization for the entire study domain $\mathcal{S}$ can be simulated using posterior estimates for $\lambda(\mathbf{u})$ $\forall\mathbf{u}:\mathbf{u}\in\mathcal{S}$. These results can then be used to characterize posterior predictive derived quantities.
4 Simulation Study
We demonstrate the multi-stage MCMC approach using a simulation study in a compact window setting. The data were first generated from an IPP with an intensity function that is log-linear in terms of three known parameters ($\beta_{0}=5$, $\beta_{1}$ = 1, $\beta_{2}$ = 2) across the entire study domain (Figure 1a). This resulted in the generation of $N=825$ total events of which $n=474$ were located within the compact observation windows (Figure 1b).
Figure 1: (a) Intensity heat map for simulated data. (b) Realization of simulated points. Points located within the compact windows (red; $n$ = 474) were used to fit the model.
Only the events located within the compact observation windows were used to fit the model with the windowed complete likelihood (3). We fit the simulated data using the first-stage sampling strategies described in section 3.2 and a conventional single-stage MCMC algorithm for comparison on a 24-core machine with 3.68 GHz processors and 192 GB of RAM. Figure 2 compares the posterior samples for $\beta_{0}$ and $\boldsymbol{\beta}$ obtained using the single-stage and the multi-stage MCMC algorithms. It is evident that the marginal posterior distributions for $\beta_{0}$ and $\boldsymbol{\beta}$ are very similar among implementation approaches. Table 2 compares the seconds per effective sample size between different methods for each parameter. The time recorded for the single-stage method does not include a priori parameter tuning, while the remaining multi-stage methods did not require tuning. Both the non-Bayesian first-stage methods showed an order of magnitude improvement. This is due to an improvement in both runtime and MCMC effective sample size. Figure 3 shows the posterior predictive distribution for $N$ obtained using the posterior samples from the GLM-E multi-stage method.
Figure 2: Comparison of marginal posterior 95% credible intervals for the single-stage method and various multi-stage methods. The true values of the parameters are shown in dashed gray.
$\beta_{0}$ $\beta_{1}$ $\beta_{2}$
Single-stage 0.59 0.34 0.26
PG 0.84 0.64 0.46
B-GLM 0.32 0.23 0.18
GLM-E 0.02 0.02 0.01
GLM-A 0.01 0.01 0.01
Table 2: Seconds per effective sample for each coefficient and method. Lower values indicate more efficient sampling.
Figure 3: Posterior predictive distribution for $N$ using posterior realizations obtained using the GLM-E method. The true value of NN is marked in dashed red.
5 Application: Harbor Seal Pups in Johns Hopkins Inlet
Johns Hopkins Inlet (Tsalxaan Niyaadé Wool’éex’i Yé; Figure 4) in Glacier Bay National Park is one of several tidewater glacial fjords in southeastern and south central Alaska that collectively host some of the largest seasonal aggregations of harbor seals (Phoca vitulina) in the world (Jansen et al., 2015; Womble et al., 2020). Harbor seals use the icebergs in these tidewater glacial fjords as haul out sites for resting, pupping, molting, and avoiding predators (Womble et al., 2020). Harbor seal abundance and space use in Johns Hopkins Inlet are of particular interest to management at Glacier Bay National Park because they help inform policy that aims to minimize potential seal disturbance while also allowing park visitors access to tidewater glaciers, one of the founding mandates of the park (National Park Service, 2010, 2025).
Harbor seal locations in Johns Hopkins Inlet were observed using aerial imagery surveys in June and August, the height of pupping and molting, respectively (Womble et al., 2020, 2021). Non-overlapping photos were taken along flight transects to avoid double-counting seals and provide systematic sampling of the entire inlet. Each image covered approximately 80 m $\times$ 120 m of surface area. We refer to the corresponding geo-referenced region as an image footprint hereafter. After imagery data were collected, a trained observer reviewed each digital image and annotated the spatial locations of adult and pup seals hauled out on ice. Further details on the survey methods and data processing are presented in Womble et al. (2020).
Previous studies analyzing these data have focused on modeling harbor seal abundance for the entire inlet or using coarse spatial grids. For example, Womble et al. (2020) incorporated the aerial imagery survey data along with other sources of data to obtain estimates and uncertainty quantification for harbor seal abundance. Womble et al. (2021) fit a multivariate conditional autoregressive (MCAR) model to summarize spatially explicit abundance on a 200 m $\times$ 300 m grid ($L=434$ total grid cells). Jansen et al. (2015) used harbor seal data from a similar aerial survey collected in 2002 to fit a generalized additive model (GAM) and summarized spatially explicit abundance on a grid of 400 m $\times$ 400 m cells ($L=779$ total grid cells). Previous studies on spatially explicit abundance of harbor seals in Johns Hopkins Inlet have not explored modeling the locations as a spatial point process directly. Modeling these data as an SPP provides unique advantages because it allows for informed simulation of potential seal locations outside the image footprints using an established statistical framework. Furthermore, the multi-stage algorithm makes fitting the model on a fine spatial grid more computationally flexible and feasible, especially with the availability of distributed computing resources.
We fit a Bayesian SPP model to harbor seal pup haul out location data from 21 June 2007 ($n=452$; Figure 5) using the aforementioned multi-stage MCMC approaches. We modeled harbor seal pups because they are the most vulnerable segment of the population and are especially susceptible to disturbance from vessels (Mathews et al., 2016). Furthermore, pups are hauled out on ice and observable with high probability at the time the survey was conducted (i.e., in the month of June and in the early afternoon) based on secondary analysis of wet-dry tag data of harbor seal pups in Disenchantment Bay, Alaska. Our study domain spanned from the Johns Hopkins Glacier terminus to Jaw Point, which is approximately 9km long and 1.5-2km wide (Figure 4a). The image footprints for this survey cover approximately half of the study domain. We assume perfect detection of hauled out pups within the image footprints and no detection outside the image footprints, giving rise to compact observation windows. Additionally, due to the systematic surveying within the study domain, we assume the unobserved areas in the inlet to be ignorable missing data (i.e., the covariate values and relationship between covariates and pup locations do not differ significantly between observed and unobserved areas). Using the pup locations as point events, we fit the SPP model using the windowed complete likelihood in (3).
Figure 4: (A) A Sentinel-2 satellite image of Johns Hopkins Inlet in 2018. Each white rectangle represents an image footprint from the 21 June 2007 survey. (B) and (C) show the location of Glacier Bay National Park and Johns Hopkins Inlet, respectively, in red.
Figure 5: Observed harbor seal pup haul out locations (red) on 21 June 2007 ($n=452$). The large polygon represents the survey domain and each gray rectangle represents an image footprint.
We fit the model using three covariates: ice proportion, bathymetry, and distance to glacier terminus (Appendix B). The ice data were represented as iceberg outlines within the image footprints and summarized as ice proportions to allow for spatial prediction outside the image footprints (Appendix C). Each covariate was summarized on the same raster grid with a spatial resolution of approximately 10 m $\times$ 10 m ($L=304,945$ total grid cells). The grid centers were used as quadrature points throughout the analysis.
5.1 Neural Network Basis Expansion
To increase model flexibility and allow for more complex relationships between covariates and intensity, we constructed a basis representation of $\mathbf{X}$, the $p$ $\times$ $L$ design matrix consisting of the covariate values at each grid center. Of the basis functions typically used to account for spatial heterogeneity (e.g., fixed-rank basis functions and radial basis functions; Hefley et al., 2017; Wikle and Zammit-Mangion, 2023), we found that a single-hidden layer feedforward neural network (SLFNN) was reasonably able to capture the relationship between the covariates and pup space use. Additionally, we found that the Extreme Learning Machine (ELM) scheme (Huang et al., 2006) was sufficient for constructing the neural network (NN) weights and resulted in efficient computation in the multi-stage algorithm.
To incorporate the NN basis expansion in the multi-stage algorithm, we first trained the NN in the first stage when fitting the logistic regression approximation. Using our samples generated from the Berman-Turner device $\Big{\{}\big{(}\mathbf{x}(\mathbf{s}_{i}),y(\mathbf{s}_{i})\big{)}\Big{\}}_{i=1}^{\tilde{n}}$, we modeled the relationship between $\mathbf{x}(\mathbf{s}_{i})$ and $\tilde{y}_{i}=\text{logit}(p_{i})$ as a linear relationship using $q$ hidden nodes
$\sum_{j=1}^{q}\tilde{\beta}_{j}g(\boldsymbol{\omega}_{j}\cdot\mathbf{x}_{i}):=\tilde{y}_{i},$ (20)
where $g(\cdot)$ represents a user-specified activation function, $\boldsymbol{\omega}_{j}$ is a $p\times 1$ vector of weights, and $\tilde{\beta}_{j}$ represents the coefficient associated with the $j$th hidden node. We omitted a bias term in (20) because we found that the weights were sufficient for constructing the basis expansion. The quantity in (20) can be summarized compactly as
$\mathbf{W}\tilde{\boldsymbol{\beta}}=\tilde{\mathbf{y}}_{i},$ (21)
where
$\mathbf{W}=\begin{bmatrix}g(\boldsymbol{\omega}_{1}\cdot\mathbf{x}_{1})&\cdots&g(\boldsymbol{\omega}_{q}\cdot\mathbf{x}_{1})\\ \vdots&\cdots&\vdots\\ g(\boldsymbol{\omega}_{1}\cdot\mathbf{x}_{\tilde{n}})&\cdots&g(\boldsymbol{\omega}_{q}\cdot\mathbf{x}_{\tilde{n}})\end{bmatrix}_{\tilde{n}\times q}.$ (22)
The matrix $\mathbf{W}$ is referred to as the hidden layer output matrix. Instead of learning the weights, $\boldsymbol{\omega}_{j}$ for $j=1,\dots,q$ using gradient-based methods, the ELM scheme samples random weights from a user-specified distribution for efficient computation Huang et al. (2006). After we computed $\mathbf{W}$ using the randomly generated weights, we used it as a basis expansion for $\mathbf{X}$ (i.e., each hidden node represents a basis vector), which then served as a new design matrix for the first-stage logistic regression problem. Aside from replacing $\mathbf{X}$ with $\mathbf{W}$, the intermediate and second stages in the multi-stage algorithm remain unchanged.
For the 21 June 2007 harbor seal pup data, we used the Gaussian Error Linear Units (GELU) activation function (Hendrycks and Gimpel, 2016) defined as $g(\mathbf{x})=\mathbf{x}\Phi(\mathbf{x})$, where $\Phi(x)$ is a standard Gaussian cumulative distribution function, and generated 100 hidden layer output matrices for $q=5$ and weights sampled from a standard Gaussian distribution. We selected the $\mathbf{W}$ matrix with the optimal AIC based on (20) for the basis expansion. We then used the GLM-E multi-stage algorithm to obtain posterior realizations for $\tilde{\boldsymbol{\beta}}$, which were used to simulate points across the entire study domain using the method described in section 3.3.2. The simulated points were summarized in a grid of posterior mean abundance (Figure 6). We validated our model using posterior predictive model-checking and did not observe large discrepancies between the simulated and observed data (Appendix D).
Figure 6: Spatially explicit posterior mean abundance of simulated harbor seal pup haul out locations on 21 June 2007.
Figure 7: Posterior predictive distribution for total abundance of hauled out seal pups NN obtained using the GLM-E posterior realizations.
When compared to the observed point pattern in Figure 5, Figure 6 indicates that greater abundance values were simulated in areas of high observed pup concentrations in addition to surrounding areas that were originally unobserved. Areas of greater simulated abundance values are also consistent with areas of higher ice proportions and more negative bathymetry values (i.e., areas of deeper water), supporting previous studies on harbor seal behavior in Southeast Alaska (Montgomery et al., 2007; Kaluzienski et al., 2023). Additionally, although we did not include this as a covariate, greater abundance values were simulated in areas where eddies typically form in the inlet, which is consistent with findings that harbor seals prefer relatively slow-moving icebergs during the pupping season (Kaluzienski et al., 2025). The posterior predictive distribution for total abundance $N$ was generated using the procedure in 3.3.1 (Figure 7). The posterior mean and standard deviation of NN were 835.2 and 27.7, respectively.
6 Discussion
Recent innovation in technology and survey techniques have led to significant improvements in the availability and spatial resolution of ecological datasets, making the development of computationally-efficient and scalable spatial point process models especially crucial. Many useful methods have been developed to facilitate flexible and efficient SPP modeling (Baddeley et al., 2010; Warton and Shepherd, 2010; Aarts et al., 2012; Fithian and Hastie, 2013) for conditional IPP likelihoods as in (2); however, fitting a Bayesian IPP model using the complete likelihood in (1), which accounts for the stochasticity associated with $n$, using conventional MCMC algorithms has remained computationally intensive.
More recently, recursive Bayes methods for fitting ecological statistical models have gained traction and have been shown to substantially improve computational efficiency (Hooten et al., 2021, 2023, 2024; Leach et al., 2022; McCaslin et al., 2021). We showed that the complete likelihood for an IPP with spatial covariates can be similarly fit in stages where the coefficients are first estimated using the likelihood conditioned on $n$, then corrected with the Poisson likelihood for $n$.
Fitting the IPP model in stages provides many benefits including increased flexibility in the first stage. We showed that fitting the first stage simplifies to fitting a model using the conditional likelihood (2), which is well approximated using logistic regression. This approximation provides various ways to fit the first stage, including user-friendly GLM R packages, making fitting Bayesian SPP models more accessible to a broader audience. Additionally, the multi-stage methods we presented do not require algorithm supervision and result in improved mixing for MCMC chains when compared to a single-stage MCMC algorithm, especially for the non-Bayesian first-stage approaches, GLM-E and GLM-A. Furthermore, all aforementioned multi-stage methods leverage parallel computing resources to pre-compute computationally intensive values, allowing the method to scale with the number of computer cores and become more feasible to fit for a large number of quadrature points.
The fast non-Bayesian estimation for logistic regression also allows for convenient pre-training of NN basis functions using the ELM scheme. A similar strategy can be applied to training other basis functions, which can be found in Hefley et al. (2017). Furthermore, tuning the parameters for a single-stage random walk MCMC algorithm becomes cumbersome as the number of basis vectors increases. Thus, it is especially useful to implement unsupervised MCMC algorithms, such as those that utilize Gibbs sampling and/or optimization-based estimation, when using a basis expansion.
One main limitation of modeling the points as an IPP is that points are assumed conditionally independent given the spatial covariates. Basis functions can allow for complex and highly non-linear relationships between covariates and point intensity, but additional dependence among points may be necessary to reasonably capture some extreme forms of spatial heterogeneity such as clustering due to social structure. Future directions of this work include extending the multi-stage procedure for self-exciting point processes such as the Hawkes process (Hawkes, 1971).
Finally, we showed how this method can be extended to the compact window setting where a posterior predictive distribution for the total abundance $N$ can be obtained. Furthermore, this method allows for intensity estimation and point simulation that lie outside the observation windows, which can be useful for downstream analyses.
7 Acknowledgments
This research was funded by the National Science Foundation Graduate Research Fellowship Program and the National Park Service. The authors thank Paul Conn, Brett McClintock, Justin Van Ee, Daniel Barreto, and Myungsoo Yoo for helpful discussions and insights. We also thank Jason Amundson and John Jansen for providing relevant data.
Appendix A
Recall that the full likelihood is given by
$\big{[}\{\mathbf{s}_{i}\},n|\beta_{0},\boldsymbol{\beta}\big{]}=\frac{\prod_{i=1}^{n}\lambda(\mathbf{s}_{i})}{n!\exp\bigl{(}\Lambda(\mathcal{S})\bigr{)}}.$
We can rewrite the full likelihood using $\zeta:=e^{\beta_{0}}$:
$\big{[}\{\mathbf{s}_{i}\},n|\beta_{0},\boldsymbol{\beta}\big{]}$ $=\frac{\prod_{i=1}^{n}\zeta\exp\{\mathbf{x}^{\prime}(\mathbf{s})\boldsymbol{\beta}\}}{n!\exp\big{\{}\zeta\Lambda(\mathcal{S})\big{\}}}$
$=\frac{1}{n!}\zeta^{n}\exp\Big{\{}-\zeta\Lambda(\mathcal{S})+\exp\{\mathbf{x}^{\prime}(\mathbf{s})\boldsymbol{\beta}\}\Big{\}}.$
Then, using a vague $\text{Gamma}(a,b)$ prior with shape-rate parameterization, we can obtain a closed-form full-conditional distribution for $\zeta$:
$\big{[}\zeta|\{\mathbf{s}_{i}\},n,\boldsymbol{\beta}\big{]}$ $\propto[\{\mathbf{s}_{i}\},n|\zeta,\boldsymbol{\beta}][\zeta]$
$\propto\zeta^{n}\exp\big{\{}-\zeta\Lambda(\mathcal{S})\big{\}}\zeta^{a-1}\exp\{-b\zeta\}$
$\propto\text{Gamma}\big{(}a+n,b+\Lambda(\mathcal{S})\big{)}.$
Recall that $\Lambda(\mathcal{S})$ in the full-conditional update can be pre-computed in the previous stage in the multi-stage MCMC algorithm. Thus, sampling from the full-conditional distribution for $\zeta$ is fast.
Appendix B
We used three covariates to fit the IPP to harbor seal pup data on 21 June 2007: bathymetry, ice proportion, and distance to glacier. For the bathymetry covariate, we downloaded a raster with cell size 1/3 arc-second ($\approx$ 10 m) from the National Oceanic and Atmospheric Administration (NOAA) Hydrographic Survey Metadata Database (HSMD). The raster values are relative to Mean Lower Low Water, with more negative values corresponding to deeper waters. We used the bathymetry raster grid to construct the remaining covariate rasters. Ice data were collected using the same aerial imagery survey described in section 5. Kaluzienski et al. (2023) conducted an ice segmentation approach to obtain iceberg outlines within each image footprint. We used the iceberg outlines to compute iceberg proportions and performed kriging to complete the ice proportion raster (Appendix C). To construct the the distance to glacier terminus raster, we computed the Euclidean distance from each grid center to the closest point on the glacier terminus.
Appendix C
The goals for utilizing the aerial ice imagery were to (1) quantify ice availability from the raw images by creating a rasterized map of ice proportions and (2) produce a map that has the same resolution and spatial extent as the other covariates being used in the model (i.e., bathymetry and distance from glacier terminus) across the entire study domain. To ensure an appropriate grid to account for the varying locations of observed ice proportions, the geometries of each image footprint in the survey date were first intersected with a pre-specified raster (Figure 8A). The resulting output was a series of grid cell geometries and their corresponding centroid coordinates.
Using these grid cell geometries, as well as a binarized version of the observed icebergs, an iterative masking method was used to find the proportion of ice within each grid cell. For each image in the survey date, the raster grid was superimposed onto the image and a logical mask was used to iterate through the grid cell and image intersections. Within a grid cell, the total number of iceberg pixels was divided by the total number of pixels in the grid cell, resulting in an ice proportion value. This process was repeated for each grid cell in an image, and for each image in the survey date (Figure 8B). Because ice was only observed within the areas of the image footprints, the calculated ice proportions and the LatticeKrig R package were used to produce an interpolated surface of ice proportions across the entire survey domain (Nychka et al., 2023). After performing kriging, a raster was produced, where ice proportion estimates were available at the same resolution as the other covariates used, and across the entire survey domain.
Figure 8: A) Pre-processed iceberg outlines for one image footprint in the 21 June 2007 survey. B) Computed ice proportions on the pre-specified covariate grid.
Appendix D
To check whether the observed data reasonably arises from the estimated inhomogeneous Poisson process, we performed a posterior predictive check based on the L-function, a function used to summarize the level of clustering in a spatial point pattern at varying distances. For each $\beta_{0}^{(k)}$ and $\boldsymbol{\beta}^{(k)}$ posterior realization obtained from the multi-stage MCMC method, we simulated points within the image footprints and computed the corresponding L-function using an isotropic edge correction. These were then compared to the L-function computed for the observed harbor seal pup locations. We found that the L-function computed from the observed pup data lie within the realm of posterior predictive L-functions (Figure 9).
Figure 9: Posterior predictive L-functions (gray) compared to the observed L-function (red).